OpenStudy (anonymous):

Can this be simplified further? -3y/5y+6 -------- y/3

OpenStudy (anonymous):

yeh a/b / c/d = a*d / b*c

OpenStudy (anonymous):

So,-9y/5y^2+6?

OpenStudy (anonymous):

the , y , will get cancelled with the one below

OpenStudy (anonymous):

um...

OpenStudy (wikiemol):

\[\frac{ \frac{ -3y}{ 5y + 6 } }{ \frac{ y }{ 3 } } = \frac{ -3y}{ 5y + 6 } * \frac{ 3 }{ y }\] dividing by a fraction is the same as multiplying by the reciprocal of that fraction.

OpenStudy (anonymous):

-3y^2/15y+18?

OpenStudy (anonymous):

then -1y/6?

OpenStudy (anonymous):

it wont be, -9y/5y^2+6 but -9y/5y^2+6y which can be simplified to -9/5y+6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw